Given are five observations for two variables, x and y. (Round your answers to two decimal places.)

xi

3 12 6 20 14

yi

55 40 55 10 15

(a)

Estimate the standard deviation of

ŷ*

when

x = 11.

(b)

Develop a 95% confidence interval for the expected value of y when

x = 11.

to

(c)

Estimate the standard deviation of an individual value of y when

x = 11.

(d)

Develop a 95% prediction interval for y when

x = 11.

to
Answer #1

a)

std error of confidence interval = s*√(1/n+(x0-x̅)2/Sxx)= 3.9158~ 3.92

b)

for 95 % confidence and 3degree of freedom critical t= 3.1820
95% confidence interval =xo -/+ t*standard error= (22.54,47.46)

c)

std error of prediction interval=s*√(1+1/n+(x0-x̅)2/Sxx)= 9.5917~ 9.59

d)

for 95 % confidence and -2degree of freedom critical t= 3.1820
95% prediction interval =xo -/+ t*standard error= (4.48,65.52)